OpenStudy (anonymous):

Suppose you invest $5000 at an annual interest rate of 6.3% compounded continuously. How much will you have in the account after 3 years? Round the solution to the nearest dollar.

jimthompson5910 (jim_thompson5910):

Use the formula A = Pe^(r*t) where A = final amount P = principle (initial amount invested) r = interest rate in decimal form t = time in years

jimthompson5910 (jim_thompson5910):

e is the constant 2.718... and it's a lot like pi = 3.14...

OpenStudy (anonymous):

okayyy thank you c:

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

what answer did you get

OpenStudy (anonymous):

the formula for these problems is A=p(1+r/n)*nt in this case it would be A=5000(1+063/1)*1*3 (exponent) A=5000(1.2) A=6000

OpenStudy (anonymous):

yeah i got $6040

jimthompson5910 (jim_thompson5910):

A = Pe^(r*t) A = 5000*e^(0.063*3) A = 6040.2047624149 A = 6040.20

jimthompson5910 (jim_thompson5910):

Round that to the nearest dollar to get $6,040 so you are correct

OpenStudy (anonymous):

okay thank you so much for ur help c:

jimthompson5910 (jim_thompson5910):

yw